### v1

# Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?

Abstract
We investigate the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction for a massive scalar to the gravitational potential, found via the non-relativistic 2-2 scattering with one and two graviton exchanges up to one loop order. Such coupling can be most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime. This coupling is qualitatively different from the minimal ones, like $\sim\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum of the graviton. We show that the insertion of one/two graviton-two scalar non-minimal vertices in the standard 2-2 scattering Feynman diagrams does not yield any contribution for the potential up to one loop, owing to the aforesaid explicit appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. In order to complement this effect, the most natural choice seems to consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant. This vertex does not contain explicitly any graviton momentum. With this, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the non-relativistic scattering amplitudes. They indeed yield a two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$. We further point out some qualitative differences of this potential with that of the standard result of $\xi=0$.
Keywords : $\xi R\phi^{2}$ interaction, perturbative quantum gravity, cosmological constant, gravitational potential
1 Introduction
There has been a lot of effort in the past and present to understand the low energy effects of perturbative quantum gravity, even though it is not renormalisable. The motivation behind this is to see its predictions at first few order of the perturbation theory, with the hope that they will be verifiable in a not too far away future. We also hope that such perturbative framework will be smoothly embedded in a complete theory of quantum gravity one day. In particular, the perturbative quantum gravity might tell us whether at all the concept of graviton as a quantum massless spin-2 field is meaningful. We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum correction to the gravitational or Newton potential, emerging from the graviton exchange of two non-relativistic particles. See also [17, 18, 19] and references therein for computation of bending of light. We also refer our reader to [20, 21, 22, 23] and references therein for computation of the corrections to Newton potential via effective field theory techniques. Moreover, see [24, 25, 26, 27, 28] for inclusion of a cosmological constant at small scales, where the background metric can still be taken to be $\eta_{\mu\nu}$. Finally, see e.g. [29, 30, 31, 32, 33] and references therein for discussions in curved spacetimes like de Sitter.
In this work we wish to investigate the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of perturbative quantum gravity for a massive scalar field, with $\xi$ being the non-minimal coupling parameter. Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self interaction [34]. What will be the contribution of this term at lowest orders in the non-minimal coupling $\xi$ and at one loop, to the correction of the Newton potential between two massive scalar particles? Note that this coupling is qualitatively different from the minimal ones. This is because for a scalar field theory with quadratic kinetic term, in the matter-graviton interaction there can be no derivative of the graviton (such as $\kappa h^{\mu\nu}T_{\mu\nu}$), whereas for the non-minimal interaction, there is no derivative of the scalar at all (6). Due to the latter, we will see that replacing any standard matter-graviton vertex by an ${\cal O}(\xi)$ one in any 2-2 scattering diagram, yields no contribution for the potential at tree or one loop level, owing to the appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude. As we need to consider the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of [3, 4]. Thus in order to find out whether at all this interaction can make any contribution to the non-relativistic potential, we must check whether there is any possibility to ‘nullify’ the effect of the aforementioned derivative or momentum of the graviton. We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three-graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2. Unlike the usual three graviton vertex generated by $2\sqrt{-g}R/\kappa^{2}$, there will be no derivative of the graviton here. Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda\kappa^{4})$. We will further emphasise qualitative differences of the potential thus found from the standard ones ($\xi=0=\Lambda$). We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solution is known to exist since long [35, 36, 37, 38]. To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ non-minimal coupling is rather less investigated from the quantum gravity perspective. Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to probe the effect of the first on quantum gravity predictions.
We also note that in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads for example in the static coordinatisation,
$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2},$ (1)
where $f=1-\Lambda r^{2}/3$. However much inside the Hubble horizon ($r\ll\sqrt{3/\Lambda}$), the spacetime can be split into a Minkowski background plus a perturbation. The Schwarzschild-de Sitter spacetime, describing the gravitational filed of a static mass inside the de Sitter universe, is given by $f(r)=1-2MG/r-\Lambda r^{2}/3$ in 1. The region with smaller $r$ values are dominated by the attraction due to the mass $M$, and the larger ones by the repulsion due to $\Lambda$. The attraction decreases monotonically with increasing $r$ whereas the repulsion increases. Accordingly, there is a radius, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g. [39]. It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, to look into short wavelength phenomena. Such decomposition in the presence of a $\Lambda$ has been done earlier in [24, 25, 26, 27, 28].
Before we go into the computations, we will summarise below the basic technical ingredients we will need. We will work with mostly positive signature of the metric in four spacetime dimensions. Our notation for symmetrisation will be : $X_{(\mu\nu)}=X_{\mu\nu}+X_{\nu\mu}$.
2 The basic ingredients
The action reads
$S=\frac{2}{\kappa^{2}}\int d^{4}x\sqrt{-g}\ \left(R-2\Lambda\right)-\frac{1}{2}\int d^{4}x\sqrt{-g}\ \left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)+m^{2}\phi^{2}+\xi R\phi^{2}\right]$ (2)
where $\kappa^{2}=32\pi G$. We wish to treat the non-minimal interaction term perturbatively. The metric is broken into background plus perturbation
$g_{\mu\nu}=\eta_{\mu\nu}+\kappa h_{\mu\nu};\qquad g^{\mu\nu}=\eta^{\mu\nu}-\kappa h^{\mu\nu}+\kappa^{2}h^{\mu}{}_{\alpha}h^{\alpha\nu}+\cdots$ (3)
Using the De Donder gauge,
$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0,$
the leading, free and linearised Einstein equation reads
$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0,$ (4)
giving
$h_{\mu\nu}=h_{\mu\nu}|_{\rm hom}-\frac{\Lambda r^{2}}{3\kappa}\eta_{\mu\nu},$ (5)
where the first term on the right hand side is the usual homogeneous massless spin-2 plane wave solution which we will use for field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$. The field quantisation with respect to $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator. We will restrict ourselves to ${\cal O}(\Lambda)$ only. We now make the following leading expansions for the scalar and the $\Lambda$ part of the action,
$S_{\phi}=-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]$
$-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]$
$+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3}),$
$S_{\Lambda}=-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2}),$ (6)
where we have used $m^{2}\gg\Lambda$, i.e. the Compton wavelength of the scalar is much small compared to the Hubble horizon, as required. The scalar and the graviton propagators read
$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}},\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right),$ (7)
where $\mu,\nu$ and $\alpha,\beta$ pairwise correspond to two graviton fields, $h_{\mu\nu}$ and $h_{\alpha\beta}$. Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4. However, this is subleading compared to the $\eta_{\mu\nu}$ term. Note also from 5 that the classical term will make contribution beyond linear order in $\Lambda$ in $S_{\Lambda}$, and hence we will take in our calculations $h_{\mu\nu}$ to be purely quantum in the cubic interaction. This term will generate a three graviton vertex of ${\cal O}(\kappa\Lambda)$ with no explicit appearance of the graviton momentum. The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g. [4],
$V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)$ $=-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right$. (8)
[…]
$\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\qquad$
with
$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right).$
We will also require the standard integral [40]
$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)}.$ (9)
Let us first see how the non-minimal term contributes to the gravitational interaction, without any $\Lambda$. The simplest process is given by the first of 1. From 6, we see that the non-minimal one graviton - two scalar vertex function reads : $i\xi\kappa(-q^{2})\eta_{\mu\nu}/2$. The $k_{2}-k^{\prime}_{2}$ vertex is the usual minimal one graviton - two scalar vertex. The Feynman amplitude reads,
$i{\cal M}=-i.i\kappa\frac{i\xi\kappa(-q^{2})\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (10)
[…]
Figure 1: The tree and the box diagrams for 2-2 scattering with graviton exchanges. Solid and wavy lines respectively denote scalar and graviton propagators. The thick circle denotes non-minimal vertex of ${\cal O}(\xi\kappa)$. The external momenta are on shell, ${k^{2}_{1}}={k^{\prime}}_{1}^{2}=-m_{1}^{2}$, ${k^{2}_{2}}={k^{\prime}}_{2}^{2}=-m_{2}^{2}$. None of the diagrams yields any non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to gravitational potential. Similar conclusion holds for the crossed box and other one loop diagrams in the absence of any $\Lambda$-three graviton vertex. See main text for detail.
3 The $\Lambda$-three graviton vertex and the gravitational potential at ${\cal O}(\xi\kappa^{4}\Lambda)$
Figure 2: The seagull diagram for 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$, with the $\Lambda$-three graviton vertex. There are three more diagrams, found by the interchange, $(k_{1},k^{\prime}_{1})\ \leftrightarrow\ (k_{2},k^{\prime}_{2})$.
Figure 3: The vacuum polarisation diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. One of the cubic graviton vertex corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8. Also as earlier, the thick circle denotes non-minimal vertex.
Figure 4: The fish diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. The cubic vertex is the $\Lambda$-three graviton vertex.
3.1 The seagull
There are total six diagrams for the seagull class, 2. Using 6 for $S_{\Lambda}$, the first of them gives the contribution ($d=4-\epsilon$)
$-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-\frac{\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
$\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
[…]
$\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$=-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
$-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,$ (11)
where $\mu$ is the mass scale for renormalisation. Likewise if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain vanishing contribution.
We next place the non-minimal vertex on the $k_{2}^{\prime}$ momentum, the second of 2. Using the on-shell and non-relativistic relations ($k_{1}^{2}={k^{\prime}_{1}}^{2}=-m_{1}^{2}$, $k_{2}^{2}={k^{\prime}_{2}}^{2}=-m_{2}^{2}$, $k_{1}\cdot k_{2}\approx-m_{1}m_{2},\,k^{\prime}_{1}\cdot k^{\prime}_{2}\approx-m_{1}m_{2}$), we have
$-i{\cal M}^{\rm 2.2}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{4}q^{2}}k_{1(\mu}k^{\prime}_{1\nu)}\eta^{\alpha\beta}\eta^{\gamma\delta}\eta^{\lambda\rho}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\mu\nu}{}_{\alpha\beta}P_{\lambda\rho\lambda^{\prime}\rho^{\prime}}P_{\gamma\delta}{}^{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\alpha^{\prime}}(l+k_{2})_{\beta^{\prime})}-l\cdot k_{2}\ \eta_{\alpha^{\prime}\beta^{\prime}}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu}{}_{\mu^{\prime}\nu^{\prime}}\eta^{\gamma\delta}P^{\nu^{\prime}\lambda^{\prime}}{}_{\gamma\delta}P_{\lambda^{\prime}}{}^{\mu^{\prime}\alpha\beta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\alpha}(l+k_{2})_{\beta)}-l\cdot k_{2}\ \eta_{\alpha\beta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}q^{2}}k_{1(\alpha}k^{\prime}_{1\beta)}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\alpha\beta}{}_{\lambda^{\prime}\rho^{\prime}}\eta^{\lambda\rho}P_{\mu\nu\lambda\rho}P^{\mu\nu\gamma\delta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\gamma}(l+k_{2})_{\delta)}-l\cdot k_{2}\ \eta_{\gamma\delta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{2}q^{2}}k_{1(\alpha}k^{\prime}_{1\beta)}P^{\alpha\beta}{}_{\mu\nu}\eta^{\lambda\rho}P_{\lambda\rho\lambda^{\prime}\rho^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\mu\nu\gamma\delta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\gamma}(l+k_{2})_{\delta)}-l\cdot k_{2}\ \eta_{\gamma\delta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})},$ (12)
which we evaluate using 9. The divergence can be absorbed in the renormalisation counterterm for the one graviton-two scalar vertex diagram embedded in 2. The part relevant to the potential (i.e., non-analytic in the transfer momentum $q^{2}$) reads
$-i{\cal M}^{\rm 2.2}\equiv-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{8\pi^{2}q^{2}}$ (13)
If we place the non-minimal vertex on the $k_{2}$ vertex, we obtain exactly the same result as above. Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result, Thus we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,
$-i{\cal M}^{\rm seagull}_{\rm tot.}\equiv-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}q^{2}}$ (14)
3.2 The vacuum polarisation
The vacuum polarisation class of diagrams are given by 3. There are two three graviton vertices involved here – one is associated with $\Lambda$, whereas the other one is the usual, 8. Using $k_{1}\cdot q\approx q^{2}/2$, $k_{1}^{\prime}\cdot q\approx-q^{2}/2$, $k_{2}\cdot q\approx-q^{2}/2$ and $k^{\prime}_{2}\cdot q\approx q^{2}/2$, we compute for the first diagram
$-i{\cal M}^{\rm 3.1}=\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0,$ (15)
where $\bar{\tau}^{\mu\nu\lambda\rho\gamma\delta}$ is the three graviton vertex of 8, modulo the factor $i\kappa/2$. Likewise, if the $k_{2}-k_{2}^{\prime}$ is non-minimal and it is connected to the $\Lambda$-three graviton vertex (i.e., the third of 3), the contribution vanishes. Similarly, we find all the contributions for this class of diagrams vanish.
3.3 The fish diagram
This class of diagrams is given by 4, and the cubic vertex here is the $\Lambda$-three graviton vertex. Let us first place the non-minimal vertex in the $k_{1}-k_{1}^{\prime}$ vertex. The two scalar-two graviton vertex is the usual (${\cal O}(\kappa)$). We thus have for the first diagram,
$-i{\cal M}^{\rm 4.1}=-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
$+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$ (16)
where we have abbreviated for the two scalar-two graviton vertex [4],
$\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}={\tau}_{2,2}^{\lambda\eta\rho\sigma}/\kappa^{2}=\left[\left\{I^{\eta\lambda\alpha\delta}I^{\rho\sigma\beta}{}_{\delta}-\frac{1}{4}\left(\eta^{\eta\lambda}I^{\rho\sigma\alpha\beta}+\eta^{\rho\sigma}I^{\eta\lambda\alpha\beta}\right)\right\}k_{2(\alpha}k^{\prime}_{2\beta)}-\frac{1}{2}\left(I^{\eta\lambda\rho\sigma}-\frac{1}{2}\eta^{\eta\lambda}\eta^{\rho\sigma}\right)(k_{2}\cdot k^{\prime}_{2}+m_{2}^{2})\right]$ (17)
A little algebra shows that 16 is vanishing. Likewise if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find vanishing contribution (the third of 4).
We next come to the case when the 2-scalar 2-graviton vertex is non-minimal, ${\cal O}(\xi\kappa^{2})$. We have for the second of 4,
$-i{\cal M}^{\rm 4.2}=\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}$
$+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}$
$-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
$\times\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}$
$=\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}$ (18)
Only the first integral contributes to the potential at leading order, which is $\sim\ln q^{2}$. Likewise by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex. Putting things together now including 14, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}$ (19)
4 The final result and discussion
The non-relativistic gravitational potential in the first Born approximation is defined as [3, 4],
$V(\vec{r})=-\frac{1}{4m_{1}m_{2}}\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}(\vec{q}^{2})$
Using the above, we obtain from 19 the leading contribution of the non-minimal coupling parameter $\xi$ to the two body gravitational potential,
$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{\mu l\hbar}{cr^{2}}\right],$ (20)
where the factor $\mu l\hbar/c$ ensures dimensional consistency, with $l$ being a length scale. In the absence of a cosmological constant, there can be no contribution from $\xi$ to the potential, at least to this order of the perturbation theory. Note that while the right hand side of the momentum space expression 19 is perfectly dimensionally consistent, the Fourier transform of the logarithm necessitates the insertion of such factor. How do we fix $l$? For example, we may set $l=R_{\rm TA,max}$, i.e. the maximum turn around radius defined below 1, as one of the natural length scales in this context, beyond which this result cannot be trusted. However, this choice is not unique and anything concrete beyond this remains elusive to us so far. Note that such ambiguity is absent for the $\xi=0$ part of the potential, due to the appearance of different powers of $G$ in different terms, fixing such term uniquely [3, 4]. (Here we do not have any term first order in $G$).
Note that the second term in the potential diverges if either of the masses is vanishing. If we understand it correctly, we do not expect such behaviour from a non-relativistic quantity. This problem originates from the absence of the scalar rest mass term in the scalar-graviton non-minimal vertex, 6. Does this imply that the $\ln\vec{q}^{2}$ term of 19 should not be related to a physical gravitational potential, or, we simply accept 20 as it is? Once again, any clear reasoning in the favour of either of this remains elusive to us so far.
Further note that the potential of 20, even though corresponds to pure short scale physics, contains two fundamental length scales (apart from $R_{\rm TA,max}$) — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$. We believe this is a novel feature encoded in a short scale physics computation.
It seems an important task to further investigate the effect of $\xi$, as well as the vertex associated with $\Lambda$ on various perturbative quantum gravity phenomenon. As we have emphasised, this coupling is qualitatively different from the minimal ones. It would thus be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertices and RG perspectives. A look into the soft theorems might also be interesting in this context. We wish to come to these issues in our future publications.
Acknowledgements : SB would like to thank Amitabha Lahiri and Parthasarathi Majumdar for useful discussions.

### v2

@@
-Can a $\xi R\phi^{2}$ non-minimal coupling contribute to the non-relativistic gravitational potential?
+
 Abstract
-We investigate the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction for a massive scalar to the gravitational potential, found via the non-relativistic 2-2 scattering with one and two graviton exchanges up to one loop order.
-Such coupling can be most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime.
-This coupling is qualitatively different from the minimal ones, like $\sim\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum of the graviton.
-We show that the insertion of one/two graviton-two scalar non-minimal vertices in the standard 2-2 scattering Feynman diagrams does not yield any contribution for the potential up to one loop, owing to the aforesaid explicit appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude.
-In order to complement this effect, the most natural choice seems to consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant.
-This vertex does not contain explicitly any graviton momentum.
-With this, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the non-relativistic scattering amplitudes.
-They indeed yield a two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$.
-We further point out some qualitative differences of this potential with that of the standard result of $\xi=0$.
+This letter investigates the contribution of the $\sqrt{-g}\xi R\phi^{2}$ non-minimal interaction to the long range gravitational potential for massive scalar fields, found from the non-relativistic limit of the 2-2 scattering amplitude with graviton exchanges.
+Such coupling is most naturally motivated from the renormalisation of a scalar field theory with quartic self interaction in a curved spacetime.
+This is qualitatively different from the minimal ones like $\sqrt{G}h^{\mu\nu}T_{\mu\nu}$, as the vertices corresponding to the former does not explicitly contain any scalar momenta, but instead explicitly contains the momentum carried by graviton line.
+For the minimal vertex, the long range gravitational potential up to one loop $({\cal O}(G),{\cal O}(G^{2}))$ was obtained earlier from the terms non-analytic in the transfer momentum, $q^{-2},\ q^{-1},\ \ln q^{2}$, yielding potentials respectively like $r^{-1}$, $r^{-2}$, $r^{-3}$.
+However owing to the aforesaid explicit appearance of graviton’s transfer momentum for the non-minimal vertices, the leading contribution in this case comes at ${\cal O}(\xi G^{2})$, and turns out to be subleading compared to even $r^{-3}$.
+In order to complement this ‘screening’ effect, we consider the three graviton vertex generated by the $\sim\Lambda\sqrt{-g}/G$ term in the action, where $\Lambda$ is the cosmological constant.
+This vertex does not contain any graviton momentum explicitly.
+With this vertex, and assuming short scale scattering much small compared to the Hubble horizon, we compute the seagull, the vacuum polarisation and the fish diagrams and obtain the 2-2 scattering amplitudes.
+The leading two body gravitational potential at ${\cal O}(\xi\Lambda G^{2})$ behaves like $r^{-1}$, even though it is much subleading compared to Newton’s potential due to the appearance of $\Lambda$.
+We also discuss the scenario where this potential dominates the aforesaid ${\cal O}(\xi G^{2})$ one.
 Keywords : $\xi R\phi^{2}$ interaction, perturbative quantum gravity, cosmological constant, gravitational potential
@@
 There has been a lot of effort in the past and present to understand the low energy effects of perturbative quantum gravity, even though it is not renormalisable.
-The motivation behind this is to see its predictions at first few order of the perturbation theory, with the hope that they will be verifiable in a not too far away future.
+The goal behind this is to see its predictions at the first few orders of the perturbation theory, with the hope that they will be verifiable in a not too far away future.
 We also hope that such perturbative framework will be smoothly embedded in a complete theory of quantum gravity one day.
 In particular, the perturbative quantum gravity might tell us whether at all the concept of graviton as a quantum massless spin-2 field is meaningful.
-We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum correction to the gravitational or Newton potential, emerging from the graviton exchange of two non-relativistic particles.
-See also [17, 18, 19] and references therein for computation of bending of light.
-We also refer our reader to [20, 21, 22, 23] and references therein for computation of the corrections to Newton potential via effective field theory techniques.
-Moreover, see [24, 25, 26, 27, 28] for inclusion of a cosmological constant at small scales, where the background metric can still be taken to be $\eta_{\mu\nu}$.
+We refer our reader to [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and references therein for various tree and one loop computations in this framework, chiefly focussing on the leading quantum corrections to the long range gravitational potential, emerging from the graviton exchanges between two non-relativistic particles.
+See [21, 22, 23] and references therein for computation of bending of light.
+We refer our reader to e.g.
+[24, 25, 26, 27] and references therein for the computations on the corrections to the Newton potential via effective field theory techniques.
+See also [28] for a technique for extracting potential from the four point amplitude in the post Minkowski expansion.
+Moreover, see [29, 30, 31, 32, 33] for inclusion of a cosmological constant at subhorizon scales, where the background metric can still be taken to be $\eta_{\mu\nu}$.
 Finally, see e.g.
-[29, 30, 31, 32, 33] and references therein for discussions in curved spacetimes like de Sitter.
-In this work we wish to investigate the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of perturbative quantum gravity for a massive scalar field, with $\xi$ being the non-minimal coupling parameter.
-Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self interaction [34].
-What will be the contribution of this term at lowest orders in the non-minimal coupling $\xi$ and at one loop, to the correction of the Newton potential between two massive scalar particles?
-Note that this coupling is qualitatively different from the minimal ones.
+[34, 35, 36, 37, 38] and references therein for discussions in curved spacetimes like the de Sitter.
+We note that the long range predictions of perturbative quantum gravity can be interpreted as an effective field theory framework making well defined physical predictions, irrespective of its ultraviolet completion.
+This serves as one of the chief motivations behind engaging into such low energy computations associated with Einstein’s or other alternative gravity theories.
+In this work we wish to probe the curvature-scalar non-minimal coupling term, $\xi\sqrt{-g}R\phi^{2}$, 2, in the context of long range gravitational potential for massive scalar fields, where $\xi$ is the non-minimal coupling parameter.
+Such coupling naturally arises in the presence of background curvature while renormalising a scalar field theory with quartic self-interaction [39, 40].
+What will be the contribution of this term at the lowest order in $\xi$, at tree and one loop level to the correction of the Newton potential?
+This coupling is qualitatively different from the minimal ones.
 This is because for a scalar field theory with quadratic kinetic term, in the matter-graviton interaction there can be no derivative of the graviton (such as $\kappa h^{\mu\nu}T_{\mu\nu}$), whereas for the non-minimal interaction, there is no derivative of the scalar at all (6).
-Due to the latter, we will see that replacing any standard matter-graviton vertex by an ${\cal O}(\xi)$ one in any 2-2 scattering diagram, yields no contribution for the potential at tree or one loop level, owing to the appearance of graviton’s transfer momentum in the numerator of the Feynman amplitude.
-As we need to consider the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of [3, 4].
-Thus in order to find out whether at all this interaction can make any contribution to the non-relativistic potential, we must check whether there is any possibility to ‘nullify’ the effect of the aforementioned derivative or momentum of the graviton.
-We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three-graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2.
+For $\xi=0$, the leading gravitational potential from 2-2 scalar scattering amplitudes up to one loop was computed earlier from the terms non-analytic in the transfer momentum carried by gravitons ($q^{-2},\ q^{-1},\ \ln q^{2}$), yielding results up to ${\cal O}(r^{-3})$, e.g.
+[4].
+For $\xi\neq 0$ however, we will see below that replacement of any minimal matter-graviton vertex by a non-minimal vertex results in the appearance of graviton’s transfer momentum in the numerator of the corresponding scattering amplitudes.
+Consequently, no long range contribution is found at tree level.
+The leading such contribution is found at ${\cal O}(\xi G^{2})$ for some one loop diagrams, and they are subleading compared to even $r^{-3}$.
+For the two scalar-one graviton and two scalar-two graviton vertices generated by the $\sqrt{-g}\xi R\phi^{2}$ term, the diagrams will have the same topology as that of $\xi=0$ [3, 4].
+Thus in order to understand whether there exists any stronger contribution, we must check if there is any way to complement this ‘screening’ effect.
+We will see that this is possible in the Einstein gravity, if we consider a cosmological constant and the corresponding ${\cal O}(\Lambda\kappa)$ three graviton vertex generated by the $4\Lambda\sqrt{-g}/\kappa^{2}$ term, 2.
 Unlike the usual three graviton vertex generated by $2\sqrt{-g}R/\kappa^{2}$, there will be no derivative of the graviton here.
-Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda\kappa^{4})$.
-We will further emphasise qualitative differences of the potential thus found from the standard ones ($\xi=0=\Lambda$).
-We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solution is known to exist since long [35, 36, 37, 38].
-To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ non-minimal coupling is rather less investigated from the quantum gravity perspective.
-Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to probe the effect of the first on quantum gravity predictions.
-We also note that in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads for example in the static coordinatisation,
-$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2},$ (1)
+Accordingly, we will see in 3 that the seagull (2), the vacuum polarisation (3) and the fish diagrams (4) will contribute to the gravitational potential, at ${\cal O}(\xi\Lambda G^{2})$, with the leading behaviour of $r^{-1}$.
+We will further emphasise qualitative differences of the potential thus found with the standard ones ($\xi=0=\Lambda$).
+We also recall that for a conformally symmetric scalar field ($m^{2}=0,\ \xi=1/6$), static spherically symmetric black hole solutions are known to exist since long [41, 42, 43, 44].
+To the best of our knowledge, the $\xi\sqrt{-g}R\phi^{2}$ coupling, even though qualitatively different from standard minimal ones, is relatively less investigated from quantum gravity perspective.
+We refer our reader for example, to [45, 46, 47, 48] for discussion on renormalisation in curved spacetime, asymptotic safety and quantum gravity, and Higgs inflation in the presence of such coupling.
+Due to its aforementioned qualitatively different nature from the standard matter-graviton coupling, it seems to be an important job to further probe the effect of the first on quantum gravity predictions.
+Now, in general the presence of a $\Lambda$ prevents us from doing quantum field theory in a flat background, and instead one should do it in a de Sitter background, which reads, for example, in the static coordinatisation,
+$ds^{2}=-fdt^{2}+f^{-1}dr^{2}+r^{2}d\Omega^{2}$ (1)
 where $f=1-\Lambda r^{2}/3$.
@@
 The region with smaller $r$ values are dominated by the attraction due to the mass $M$, and the larger ones by the repulsion due to $\Lambda$.
-The attraction decreases monotonically with increasing $r$ whereas the repulsion increases.
-Accordingly, there is a radius, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g.
+The attraction decreases monotonically with increasing $r$, whereas the repulsion increases.
+Accordingly, there is a scale, $R_{\rm TA,max}=(3MG/\Lambda)^{1/3},$ known as the maximum turn around radius, around which the attraction and repulsion balance each other, e.g.
 [49].
-It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, to look into short wavelength phenomena.
+It is clear that we may do a decomposition of flat spacetime plus perturbation inside this radius, essentially to look into short wavelength phenomena.
 Such decomposition in the presence of a $\Lambda$ has been done earlier in [29, 30, 31, 32, 33].
+One may reasonably anticipate that these short scale effects will be tiny while compared to the $\Lambda=0$ cases hitherto investigated.
+Nevertheless, we believe probing such effects are important, in order to further understand the nature of the matter-gravity interaction.
+We also note that in the primordial inflationary scenario and in the infrared regime, one can encounter non-perturbative secular large logarithms while computing the quantum corrections to the two body gravitational potential, e.g.
+[36] and references therein.
+Understanding the role of the $\xi\sqrt{-g}R\phi^{2}$ coupling in the inflationary and large scale context can be an important task, which we reserve for the future.
+We wish to emphasise that the phrases short scale/short wavelength appearing anywhere in this letter is to be understood as the subhorizon/subHubble scale.
+Even though the potential is long range, it is valid only up to the scale of the maximum turn around radius of a structure.
+Beyond this scale the expansion of the universe dominates, and the Hubble flow will essentially invalidate the notion of a potential as an explicit time independent non-relativistic quantity.
+In this regard, we also note that while the maximum turn around radius mentioned above can be thought of as a local test of cosmological constant/dark energy [49], the other effects of $\Lambda$ on the gravitational physics inside the maximum turn around radius are expected to be small compared to that of the mass of the gravitating object.
+$\Lambda$ dominates at very large scales beyond the maximum turn around radius, and contributes to phenomena like the gravitational redshift, or correlations originating from the very early universe, etc..
+Nevertheless, with increasing measurement facilities such as the event horizon telescope, one can hope to probe the effect of $\Lambda$ in short scale physics (like the solar system) as well, in the not too far away future, like the perihelion precession, bending of light and the Shapiro delay.
+This might give us new insight about the nature of the dark energy.
+Thus attempts to understand such phenomena’s perturbative quantum gravity corrections are also well motivated.
 Before we go into the computations, we will summarise below the basic technical ingredients we will need.
-We will work with mostly positive signature of the metric in four spacetime dimensions.
+We will work with the mostly positive signature of the metric in four spacetime dimensions.
 Our notation for symmetrisation will be : $X_{(\mu\nu)}=X_{\mu\nu}+X_{\nu\mu}$.
 2 The basic ingredients
-The action reads
+Our action reads
 $S=\frac{2}{\kappa^{2}}\int d^{4}x\sqrt{-g}\ \left(R-2\Lambda\right)-\frac{1}{2}\int d^{4}x\sqrt{-g}\ \left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)+m^{2}\phi^{2}+\xi R\phi^{2}\right]$ (2)
@@
 Using the De Donder gauge,
-$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0,$
+$\partial_{\mu}\left(h^{\mu}{}_{\nu}-\frac{1}{2}\delta^{\mu}_{\nu}h\right)=0$
 the leading, free and linearised Einstein equation reads
-$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0,$ (4)
+$\partial^{2}\left(h_{\mu\nu}-\frac{1}{2}h\eta_{\mu\nu}\right)-\frac{2\Lambda}{\kappa}\eta_{\mu\nu}=0$ (4)
 giving
 $h_{\mu\nu}=h_{\mu\nu}|_{\rm hom}-\frac{\Lambda r^{2}}{3\kappa}\eta_{\mu\nu},$ (5)
-where the first term on the right hand side is the usual homogeneous massless spin-2 plane wave solution which we will use for field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$.
-The field quantisation with respect to $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator.
-We will restrict ourselves to ${\cal O}(\Lambda)$ only.
+where the first term on the right-hand side is the usual homogeneous massless spin-2 plane wave solution, which will give the field quantisation, whereas the second term corresponds to the linearised, classical de Sitter perturbation yielding the background curvature, $R=4\Lambda$.
+The quantisation for $h_{\mu\nu}|_{\rm hom}$ yields the graviton propagator.
+The ground state for the theory will be taken to be the standard Poincarré invariant Minkowski vacuum, subject to the flat background, 3.
+Note that in the presence of $\Lambda$, this cannot be possible at large scales, due to particle creation and the subsequent instability of the initial vacuum state.
+This happens if we are concerned with length scales comparable to or higher than the Hubble horizon.
+Such large scale, infrared effects are essentially non-perturbative and non-equilibrium, so that the standard in-out $S$-matrix formalism of quantum field theory which we will be using here, fails.
+Scattering has no well defined meaning in such dynamical backgrounds and one needs to do the in-in formalism to compute expectation values or correlation functions, e.g.
+[38] and references therein.
+This problem is not relevant to our present scenario, for as we have stated in the preceding section, we are concerned here only up to the length scale of a structure (like a galaxy) defined by its maximum turn around radius.
+This length scale is much subhorizon and hence we may safely ignore the aforementioned large scale IR effects for our present purpose.
+This means that the $\Lambda$ term appearing in 5 is much small compared to unity.
+We will assert this scenario in our calculations in the following by taking the Compton wavelength of the scattering particles to be much small compared to the horizon length.
+Also, since the current observed value of $\Lambda$ is tiny, ${\cal O}(10^{-52}\ {\rm m}^{-2})$, we will restrict our computations to ${\cal O}(\Lambda)$ only.
 We now make the following leading expansions for the scalar and the $\Lambda$ part of the action,
-$S_{\phi}=-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]$
-$-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]$
-$+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3}),$
-$S_{\Lambda}=-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2}),$ (6)
-where we have used $m^{2}\gg\Lambda$, i.e. the Compton wavelength of the scalar is much small compared to the Hubble horizon, as required. The scalar and the graviton propagators read
-$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}},\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right),$ (7)
+$\begin{split}S_{\phi}=&-\frac{1}{2}\int d^{4}x\left[(\partial\phi)^{2}+m^{2}\phi^{2}\right]+\frac{\kappa}{2}\int d^{4}xh^{\mu\nu}\left[(\partial_{\mu}\phi)(\partial_{\nu}\phi)-\frac{1}{2}\eta_{\mu\nu}\left((\partial\phi)^{2}+m^{2}\phi^{2}\right)\right]\\ &-\frac{\kappa^{2}}{2}\int d^{4}x\left[\left(h^{\mu}{}_{\alpha}h^{\alpha\nu}-\frac{1}{2}hh^{\mu\nu}+\frac{h^{2}}{8}\eta^{\mu\nu}-\frac{1}{4}h_{\alpha\beta}h^{\alpha\beta}\eta^{\mu\nu}\right)(\partial_{\mu}\phi)(\partial_{\nu}\phi)+\frac{m^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)\phi^{2}\right]\\ &+\frac{\xi}{2}\int d^{4}x\left[\kappa\phi^{2}\partial^{2}h-\kappa^{2}\phi^{2}\left\{-\frac{1}{4}h\partial^{2}h-\frac{1}{2}(\partial^{\lambda}h^{\mu\alpha})(\partial_{\alpha}h_{\mu\lambda})+h^{\mu\nu}\partial^{2}h_{\mu\nu}+\frac{3}{4}(\partial^{\lambda}h^{\mu\nu})(\partial_{\lambda}h_{\mu\nu})\right\}\right]+\ {\cal O}(\kappa^{3},\xi\kappa^{3})\\ S_{\Lambda}=&-\frac{4\Lambda}{\kappa^{2}}\int d^{d}x\left[1+\frac{\kappa h}{2}+\frac{\kappa^{2}}{4}\left(\frac{h^{2}}{2}-h_{\mu\nu}h^{\mu\nu}\right)+\frac{\kappa^{3}}{2}\left(\frac{1}{3}h_{\mu\nu}h^{\nu\lambda}h_{\lambda}{}^{\mu}-\frac{1}{4}hh_{\mu\nu}h^{\mu\nu}+\frac{h^{3}}{24}\right)\right]+\ {\cal O}(\Lambda\kappa^{2})\end{split}$ (6)
+where we have used $m^{2}\gg\Lambda$, i.e., the Compton wavelength of the scalar is much small compared to the Hubble horizon, as stated above.
+Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4.
+However, this is subleading compared to the $\eta_{\mu\nu}$ term, and hence has been ignored.
+Also as 5 suggests, both quantum $h_{\mu\nu}|_{\rm hom}$ and the classical $\Lambda r^{2}$ terms make contributions in the above.
+However, we will only be concerned about the quantum effects generated by $h_{\mu\nu}|_{\rm hom}$, as the classical terms cannot contribute in forming a vertex or in carrying momenta.
+Moreover in $S_{\Lambda}$, the classical $\Lambda r^{2}$ term will make contributions beyond linear order in $\Lambda$, hence are further subleading in our subhorizon scenario.
+Thus we have a three graviton vertex of ${\cal O}(\kappa\Lambda)$ from $S_{\Lambda}$.
+There is no explicit appearance of the graviton momentum here, for it contains no derivative of the graviton.
+This vertex will serve the key role in our following computations.
+The scalar and the graviton propagators read
+$\Delta(k)=-\frac{1}{k^{2}+m^{2}};\qquad\Delta_{\mu\nu\alpha\beta}(k)=-\frac{P_{\mu\nu\alpha\beta}}{k^{2}};\qquad P_{\mu\nu\alpha\beta}=\frac{1}{2}\left(\eta_{\mu\alpha}\eta_{\nu\beta}+\eta_{\mu\beta}\eta_{\nu\alpha}-\eta_{\mu\nu}\eta_{\alpha\beta}\right)$ (7)
 where $\mu,\nu$ and $\alpha,\beta$ pairwise correspond to two graviton fields, $h_{\mu\nu}$ and $h_{\alpha\beta}$.
-Note in 6 that there is a term quadratic in $h_{\mu\nu}$ in $S_{\Lambda}$, which could also make a contribution proportional to $\Lambda h_{\mu\nu}$ in 4.
-However, this is subleading compared to the $\eta_{\mu\nu}$ term.
-Note also from 5 that the classical term will make contribution beyond linear order in $\Lambda$ in $S_{\Lambda}$, and hence we will take in our calculations $h_{\mu\nu}$ to be purely quantum in the cubic interaction.
-This term will generate a three graviton vertex of ${\cal O}(\kappa\Lambda)$ with no explicit appearance of the graviton momentum.
-The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g.
-[4],
-$V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)$ $=-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right$.
-(8)
-[…]
-[…]
-[…]
-[…]
-[…]
-$\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\qquad$
-with
-$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right).$
-We will also require the standard integral [40]
-$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)}.$ (9)
+The usual three graviton vertex originating from the $\sqrt{-g}R/2\kappa^{2}$ term (with momenta $l$, $q$, lql-q along the three lines) reads, e.g. [4],
+$\begin{split}V^{(3)\mu\nu}_{\alpha\beta\gamma\delta}(\kappa;l,q)=&-\frac{i\kappa}{2}\left[P_{\alpha\beta\gamma\delta}\left(l^{\mu}l^{\nu}+(l-q)^{\mu}(l-q)^{\nu}+q^{\mu}q^{\nu}-\frac{3}{2}\eta^{\mu\nu}q^{2}\right)\right.\\ &\left.+2q_{\lambda}q_{\sigma}\left(I_{\alpha\beta}{}^{\sigma\lambda}I_{\gamma\delta}{}^{\mu\nu}+I_{\gamma\delta}{}^{\sigma\lambda}I_{\alpha\beta}{}^{\mu\nu}-I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta}{}^{\nu\lambda}-I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta}{}^{\nu\lambda}\right)\right.\\ &\left.+\left\{q_{\lambda}q^{\mu}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\nu\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\nu\lambda}\right)+q_{\lambda}q^{\nu}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\mu\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\mu\lambda}\right)-q^{2}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\mu\nu}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\mu\nu}\right)\right.\right.\\ &\left.\left.-\eta^{\mu\nu}q_{\sigma}q_{\lambda}\left(\eta_{\alpha\beta}I_{\gamma\delta}{}^{\sigma\lambda}+\eta_{\gamma\delta}I_{\alpha\beta}{}^{\sigma\lambda}\right)\right\}\right.\\ &\left.+\left\{-2q_{\lambda}\left(I_{\alpha\beta}{}^{\lambda\sigma}I_{\gamma\delta\sigma}{}^{\nu}(l-q)^{\mu}+I_{\alpha\beta}{}^{\lambda\sigma}I_{\gamma\delta\sigma}{}^{\mu}(l-q)^{\nu}+I_{\gamma\delta}{}^{\lambda\sigma}I_{\alpha\beta\sigma}{}^{\nu}l^{\mu}+I_{\gamma\delta}{}^{\lambda\sigma}I_{\alpha\beta\sigma}{}^{\mu}l^{\nu}\right)\right.\right.\\ &\left.\left.+q^{2}\left(I_{\alpha\beta\sigma}{}^{\mu}I_{\gamma\delta}{}^{\nu\sigma}+I_{\gamma\delta\sigma}{}^{\mu}I_{\alpha\beta}{}^{\nu\sigma}\right)+\eta^{\mu\nu}q_{\sigma}q_{\lambda}\left(I_{\alpha\beta}{}^{\lambda\rho}I_{\gamma\delta\rho}{}^{\sigma}+I_{\gamma\delta}{}^{\lambda\rho}I_{\alpha\beta\rho}{}^{\sigma}\right)\right\}\right.\\ &\left.+\left\{\left(l^{2}+(l-q)^{2}\right)\left(I_{\alpha\beta}{}^{\mu\sigma}I_{\gamma\delta\sigma}{}^{\nu}+I_{\gamma\delta}{}^{\mu\sigma}I_{\alpha\beta\sigma}{}^{\nu}-\frac{1}{2}\eta^{\mu\nu}{\cal P}_{\alpha\beta\gamma\delta}\right)-\left(I_{\gamma\delta}{}^{\mu\nu}\eta_{\alpha\beta}l^{2}+I_{\alpha\beta}{}^{\mu\nu}\eta_{\gamma\delta}(l-q)^{2}\right)\right\}\right],\end{split}$ (8)
+where
+$I_{\mu\nu\lambda\rho}=\frac{1}{2}\left(\eta_{\mu\lambda}\eta_{\nu\rho}+\eta_{\mu\rho}\eta_{\nu\lambda}\right)$
+We will use the standard integral [50]
+$I^{r}_{s}=\int\frac{d^{d}k}{(2\pi)^{d}}\frac{(k^{2})^{r}}{(k^{2}-a^{2})^{s}}=\frac{i(-1)^{r+s}}{(4\pi)^{d/2}}\frac{1}{(a^{2})^{s-r-d/2}}\frac{\Gamma(r+d/2)\Gamma(s-r-d/2)}{\Gamma(d/2)\Gamma(s)},$ (9)
+as well as the Fourier transforms
+$\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\frac{e^{i\vec{q}\cdot\vec{r}}}{\vec{q}^{2}}=\frac{1}{4\pi r},\qquad\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}\frac{e^{i\vec{q}\cdot\vec{r}}}{|\vec{q}|}=\frac{1}{2\pi^{2}r^{2}},\qquad\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}\ln\vec{q}^{2}=-\frac{1}{2\pi r^{3}}$ (10)
+in the following calculations.
+Also, we will take the dummy virtual momentum named $l^{\mu}$ appearing below to run from -$\infty$ to $\infty$.
+The amplitudes will be evaluated in the usual relativistic framework, whereas the non-relativistic limit (such as $k_{1}\cdot k_{2}\approx-m_{1}m_{2}$) will be taken after the evaluation.
 Let us first see how the non-minimal term contributes to the gravitational interaction, without any $\Lambda$.
-The simplest process is given by the first of 1.
+The tree level process is given by the first of 1.
 From 6, we see that the non-minimal one graviton-two scalar vertex function reads: $i\xi\kappa(-q^{2})\eta_{\mu\nu}/2$.
-The $k_{2}-k^{\prime}_{2}$ vertex is the usual minimal one graviton - two scalar vertex.
+The $k_{2}-k^{\prime}_{2}$ vertex is taken to be the usual minimal one graviton-two scalar vertex.
 The Feynman amplitude reads,
-$i{\cal M}=-i.i\kappa\frac{i\xi\kappa(-q^{2})\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (10)
-[…]
+$i{\cal M}=-i.i\kappa\frac{i(-q^{2})\xi\kappa\eta_{\mu\nu}P^{\mu\nu\alpha\beta}\left(k_{1(\alpha}k^{\prime}_{1\beta)}-\eta_{\alpha\beta}(k_{1}\cdot k^{\prime}_{1}+m_{1}^{2})\right)}{2^{2}q^{2}}$ (11)
+where $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$.
+The long range gravitational potential is defined in the non-relativistic limit by the Fourier transform of ${\cal M}/4m_{1}m_{2}$ in the first Born approximation, originating from the functions non-analytic in $q^{2}$.
+However, the potential corresponding 11 is proportional to $\delta^{3}(\vec{r})$, i.e., not a long range one we are interested in.
+It can be checked explicitly that analogous things occurs for the box diagram of 1.
+As a further explicit example, let us consider the triangle diagram generated by pinching the $k_{1}-l$-line of the box diagram.
+This will generate a two scalar-two graviton vertex in the upper half of the diagram.
+Let us first imagine that this vertex is non-minimal, produced by 6.
+It is not difficult to see that this diagram generates a momentum integral of the kind
+$\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{l\cdot(l-q)\ k_{2}\cdot(l-k^{\prime}_{2})\ k^{\prime}_{2}\cdot(l-k^{\prime}_{2})}{l^{2}(l-q)^{2}\left((l-k^{\prime}_{2})^{2}+m_{2}^{2}\right)},$ (12)
+which gives a leading potential $\sim G^{2}\xi m_{2}^{2}/m_{1}r^{4}$ in the non-relativistic limit.
+Note the appearance of the mass of the scalar in the denominator.
+Even though long range, this potential is much subleading and we wish to look for leading results than this.
+Similar result is found if instead we pinch the $(l+k_{2})$-line of the second of 1.
+If on the other hand, we take either the $k_{2}$ or the $k^{\prime}_{2}$ vertex to be non-minimal, we encounter integrals like
+$\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{l^{2}\ k_{2}\cdot(l+k_{2})}{l^{2}(l-q)^{2}\left((l+k_{2})^{2}+m_{2}^{2}\right)}\quad{\rm Or,}\quad\sim\int\frac{d^{d}l}{(2\pi)^{d}}\frac{(l-q)^{2}\ k^{\prime}_{2}\cdot(l+k_{2})}{l^{2}(l-q)^{2}\left((l+k_{2})^{2}+m_{2}^{2}\right)},$
+both of which are independent of $q$ and hence makes no contribution to the long range potential.
+Likewise, for the other one loop diagrams as well, we find either they do not yield any potential at all, or their leading contribution is ${\cal O}(r^{-4})$.
+As we have already emphasised, evidently the above situation arises due to the transfer momentum carried by the graviton appearing in the numerator.
+This motivates us to bring in the $\Lambda$-three graviton vertex mentioned above, and to see whether we may obtain contribution leading compared to $r^{-4}$.
+There can be three relevant diagrams involving this vertex at one loop, which we wish to evaluate below.
+9 will be one of the chief ingredients for the following calculations.
 Figure 1: The tree and the box diagrams for 2-2 scattering with graviton exchanges.
 Solid and wavy lines respectively denote scalar and graviton propagators.
-The thick circle denotes non-minimal vertex of ${\cal O}(\xi\kappa)$.
+The thick circle denotes a non-minimal vertex of ${\cal O}(\xi\kappa)$.
 The external momenta are on shell, ${k^{2}_{1}}={k^{\prime}}_{1}^{2}=-m_{1}^{2}$, ${k^{2}_{2}}={k^{\prime}}_{2}^{2}=-m_{2}^{2}$.
-None of the diagrams yields any non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to gravitational potential.
-Similar conclusion holds for the crossed box and other one loop diagrams in the absence of any $\Lambda$-three graviton vertex.
-See main text for detail.
+Both diagrams yield no non-analytic contribution in the transfer momentum, $q=k_{1}-k^{\prime}_{1}=k^{\prime}_{2}-k_{2}$, and hence to the long range gravitational potential.
+The leading contribution behaves as $r^{-4}$ at ${\cal{O}}(\xi G^{2})$ from some other diagrams.
+See the main text for details.
 3 The $\Lambda$-three graviton vertex and the gravitational potential at ${\cal O}(\xi\kappa^{4}\Lambda)$
@@
 Figure 3: The vacuum polarisation diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$.
-One of the cubic graviton vertex corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8.
-Also as earlier, the thick circle denotes non-minimal vertex.
+One of the cubic graviton vertices corresponds to the $\Lambda$-three graviton vertex, 6, whereas the other one is the usual, 8.
+Also, as earlier, the thick circle denotes a non-minimal vertex.
 Figure 4: The fish diagram for the 2-2 scalar scattering at ${\cal O}(\xi\kappa^{4}\Lambda)$. The cubic vertex is the $\Lambda$-three graviton vertex.
@@
 There are total six diagrams for the seagull class, 2. Using 6 for $S_{\Lambda}$, the first of them gives the contribution ($d=4-\epsilon$)
-$-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
-$\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$-\frac{\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}$
-$\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-[…]
-$\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$+\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$=-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}$
-$-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,$ (11)
-where $\mu$ is the mass scale for renormalisation. Likewise if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain vanishing contribution.
+$\begin{split}&-i{\cal M}^{\rm 2.1}=\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{3}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda\rho}\eta_{\gamma\delta}P^{\lambda\rho\lambda^{\prime}\rho^{\prime}}P^{\gamma\delta\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\lambda^{\prime}}(l-k^{\prime}_{2})_{\rho^{\prime})}-\eta_{\lambda^{\prime}\rho^{\prime}}(k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\gamma^{\prime}}(l-k_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}P_{\mu\nu\alpha\beta}\eta^{\alpha\beta}\eta_{\lambda^{\prime}\mu^{\prime}}\eta_{\nu^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\mu^{\prime}\nu^{\prime}\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2}\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\mu\nu}\eta^{\mu^{\prime}\lambda^{\prime}}\eta^{\nu^{\prime}\rho^{\prime}}\eta_{\alpha\beta}P_{\mu\nu\mu^{\prime}\nu^{\prime}}P^{\alpha\beta\gamma\delta}P_{\lambda^{\prime}\rho^{\prime}}{}^{\gamma^{\prime}\delta^{\prime}}\left\{k^{\prime}_{2(\gamma^{\prime}}(l-k^{\prime}_{2})_{\delta^{\prime})}-\eta_{\gamma^{\prime}\delta^{\prime}}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\\ &\times\frac{\left\{k_{2(\gamma}(l-k_{2})_{\delta)}-\eta_{\gamma\delta}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+\mu^{4\epsilon}\Lambda\kappa^{4}\xi\int\frac{d^{d}l}{(2\pi)^{d}}q^{2}\eta^{\alpha\beta}P_{\alpha\beta\mu\nu}P^{\nu\lambda\gamma\delta}P_{\lambda}{}^{\mu\alpha^{\prime}\beta^{\prime}}\left\{k^{\prime}_{2(\gamma}(l-k^{\prime}_{2})_{\delta)}-\eta_{\gamma\delta}(k^{\prime}_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2})\right\}\frac{\left\{k_{2(\alpha^{\prime}}(l-k_{2})_{\beta^{\prime})}-\eta_{\alpha^{\prime}\beta^{\prime}}(k_{2}\cdot(l-k_{2})-m_{2}^{2})\right\}}{q^{2}l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ =&-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k^{\prime}_{2}\cdot(l-k^{\prime}_{2})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &+2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot(l-k_{2})k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-2m_{2}^{2}k_{2}\cdot(l-k_{2})+4m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}\\ &-2\Lambda\kappa^{4}\xi\mu^{4\epsilon}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2}\cdot k_{2}^{\prime}(l-k_{2}^{\prime})\cdot(l-k_{2})+k_{2}^{\prime}\cdot(l-k_{2})k_{2}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}^{\prime}\cdot(l-k_{2}^{\prime})-m_{2}^{2}k_{2}\cdot(l-k_{2})+2m_{2}^{4}}{l^{2}(l-q)^{2}((l-k_{2}^{\prime})^{2}+m_{2}^{2})}=0,\end{split}$ (13)
+where $\mu$ is the mass scale for renormalisation. Likewise, if we interchange pairwise $k_{1},k^{\prime}_{1}$ with $k_{2},k^{\prime}_{2}$, we obtain a vanishing contribution.
 We next place the non-minimal vertex on the $k_{2}^{\prime}$ momentum, the second of 2.
@@
 $-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2^{2}q^{2}}k_{1(\alpha}k^{\prime}_{1\beta)}P^{\alpha\beta}{}_{\mu\nu}\eta^{\lambda\rho}P_{\lambda\rho\lambda^{\prime}\rho^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}P^{\mu\nu\gamma\delta}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{k_{2(\gamma}(l+k_{2})_{\delta)}-l\cdot k_{2}\ \eta_{\gamma\delta}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
-$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
-$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
-$+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
-$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}$
-$=-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]$
-$-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})},$ (12)
+$\begin{split}=&-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\\ &+\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}+\frac{1}{l^{2}}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\end{split}$
+$\begin{split}=&-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi m_{1}^{2}}{4q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\left[\frac{2m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}-\frac{1}{(l+k_{2})^{2}+m_{2}^{2}}\right]-\frac{\mu^{4\epsilon}\Lambda\kappa^{4}\xi}{2q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2l\cdot\left(m_{1}m_{2}k_{1}^{\prime}+m_{1}m_{2}k_{1}-m_{1}^{2}k_{2}\right)-4m_{1}^{2}m_{2}^{2}}{l^{2}((l+k_{2})^{2}+m_{2}^{2})}\end{split}$ (14)
 which we evaluate using 9.
@@
 If we place the non-minimal vertex on the $k_{2}$ vertex, we obtain exactly the same result as above.
-Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result, Thus we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,
+Likewise, if we pairwise interchange the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices, we obtain the same result.
+Thus, we have the total contribution to the gravitational potential for the seagull class of diagrams, 2,
 $-i{\cal M}^{\rm seagull}_{\rm tot.}\equiv-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}q^{2}}$ (16)
+By 10, the above amplitude will give an $r^{-1}$ potential in the non-relativistic limit.
+Note that there is no subleading contribution to the potential from the seagull diagrams.
+This is simply because only the second and the third diagrams of 3 (along with the pairwise interchange of the $k_{1},k_{1}^{\prime}$ and $k_{2},k_{2}^{\prime}$ vertices) contribute to the above potential.
+Due to the placement of the nonminimal vertex in these diagrams we have $(l-q)^{2}$ or $l^{2}$ appearing in the numerator of the corresponding scattering amplitudes.
+This results in the factorisation of $q^{-2}$ and momentum integral independent of the transfer momentum $q$, 14.
+Accordingly, the contribution in 16 is exact, and there is no other terms (like $q^{-1}$ or $\ln q^{2}$) which can make subleading contributions to the potential.
 3.2 The vacuum polarisation
@@
 Using $k_{1}\cdot q\approx q^{2}/2$, $k_{1}^{\prime}\cdot q\approx-q^{2}/2$, $k_{2}\cdot q\approx-q^{2}/2$ and $k^{\prime}_{2}\cdot q\approx q^{2}/2$, we compute for the first diagram
-$-i{\cal M}^{\rm 3.1}=\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}$
-$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
-$+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}$
-$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
-$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
-$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}$
-$-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}$
-$\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0,$ (15)
+$\begin{split}-i{\cal M}^{\rm 3.1}=&\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{8}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}\eta^{\lambda^{\prime}\rho^{\prime}}\eta^{\gamma^{\prime}\delta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P_{\gamma^{\prime}\delta^{\prime}\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}+3\mu^{4\epsilon}\kappa^{4}\Lambda\xi(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\beta^{\prime}\gamma^{\prime}}{}_{\lambda\rho}P_{\gamma^{\prime}}{}^{\alpha^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{4}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\lambda\rho}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}-\frac{3\mu^{4\epsilon}\kappa^{4}\Lambda\xi}{2}(-q^{2})\frac{\eta^{\mu^{\prime}\nu^{\prime}}}{q^{2}}P_{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\rho}\eta_{\lambda^{\prime}\rho^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\gamma\delta}\\ &\times\left\{k_{2(\mu}k^{\prime}_{2\nu)}-\eta_{\mu\nu}(k_{2}\cdot k_{2}^{\prime}+m^{2}_{2})\right\}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\bar{\tau}^{\mu^{\prime\prime}\nu^{\prime\prime}\lambda\rho\gamma\delta}}{l^{2}(l-q)^{2}}\frac{1}{q^{2}}P_{\mu^{\prime\prime}\nu^{\prime\prime}}{}^{\mu\nu}=0\end{split}$ (17)
 where $\bar{\tau}^{\mu\nu\lambda\rho\gamma\delta}$ is the three graviton vertex of 8, modulo the factor $i\kappa/2$.
 Likewise, if the $k_{2}-k_{2}^{\prime}$ is non-minimal and it is connected to the $\Lambda$-three graviton vertex (i.e., the third of 3), the contribution vanishes.
-Similarly, we find all the contributions for this class of diagrams vanish.
+Similarly, we find that all the contributions for this class of diagrams vanish.
 3.3 The fish diagram
@@
 We thus have for the first diagram,
-$-i{\cal M}^{\rm 4.1}=-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
-$-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
-$+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$
-$+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}$ (16)
+$\begin{split}-i{\cal M}^{\rm 4.1}=&-\Lambda\mu^{4\epsilon}\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\lambda^{\prime}\rho^{\prime}}}{q^{2}}\eta^{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{\eta_{\alpha^{\prime}\beta^{\prime}}P^{\alpha^{\prime}\beta^{\prime}}{}_{\lambda\eta}\eta_{\rho^{\prime}\sigma^{\prime}}P^{\rho^{\prime}\sigma^{\prime}}{}_{\rho\sigma}}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\\ &-8\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\nu^{\prime}\lambda^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}}{}^{\mu^{\prime}}_{\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\end{split}$ (18)
+$\begin{split}&+2\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}\eta^{\mu^{\prime}\nu^{\prime}}P^{\lambda^{\prime}\rho^{\prime}}{}_{\lambda\eta}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\\ &+4\mu^{4\epsilon}\Lambda\kappa^{4}\xi(-q)^{2}\eta^{\mu\nu}\frac{P_{\mu\nu\mu^{\prime}\nu^{\prime}}}{q^{2}}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda\eta}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\rho\sigma}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}\times\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}\end{split}$
 where we have abbreviated for the two scalar-two graviton vertex [4],
 $\bar{\tau}_{2,2}^{\lambda\eta\rho\sigma}={\tau}_{2,2}^{\lambda\eta\rho\sigma}/\kappa^{2}=\left[\left\{I^{\eta\lambda\alpha\delta}I^{\rho\sigma\beta}{}_{\delta}-\frac{1}{4}\left(\eta^{\eta\lambda}I^{\rho\sigma\alpha\beta}+\eta^{\rho\sigma}I^{\eta\lambda\alpha\beta}\right)\right\}k_{2(\alpha}k^{\prime}_{2\beta)}-\frac{1}{2}\left(I^{\eta\lambda\rho\sigma}-\frac{1}{2}\eta^{\eta\lambda}\eta^{\rho\sigma}\right)(k_{2}\cdot k^{\prime}_{2}+m_{2}^{2})\right]$ (19)
-A little algebra shows that 16 is vanishing. Likewise if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find vanishing contribution (the third of 4).
+A little algebra shows that 18 is vanishing. Likewise, if the ${\cal O}(\kappa)$ non-minimal vertex is $k_{2}-k^{\prime}_{2}$, we find a vanishing contribution (the third of 4).
 We next come to the case when the 2-scalar 2-graviton vertex is non-minimal, ${\cal O}(\xi\kappa^{2})$. We have for the second of 4,
-$-i{\cal M}^{\rm 4.2}=\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
-$\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}$
-$+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
-$\times\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}$
-$-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
-$\times\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}$
-$-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}$
-$\times\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}$
-$=\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}$ (18)
+$\begin{split}-i{\cal M}^{\rm 4.2}=&\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{4}k_{1(\mu}k^{\prime}_{1\nu)}P^{\mu\nu\mu^{\prime}\nu^{\prime}}\eta_{\mu^{\prime}\nu^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\\ &\times\left\{-\frac{1}{2}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\alpha}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}}{}^{\mu\nu}\eta^{\alpha^{\prime}\beta^{\prime}}P_{\alpha^{\prime}\beta^{\prime}\mu\nu}\right\}\\ &+2\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}}{}_{\lambda^{\prime}\rho^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P^{\rho^{\prime}\gamma^{\prime}\mu\alpha}{P_{\gamma^{\prime}}{}^{\lambda^{\prime}}}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P^{\gamma^{\prime}\rho^{\prime}\mu\nu}P^{\lambda^{\prime}}{}_{\gamma^{\prime}\mu\nu}\right\}\\ &-\mu^{3\epsilon}\Lambda\kappa^{4}\xi k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\alpha}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\alpha^{\prime}\beta^{\prime}}{}^{\mu\nu}\eta^{\lambda^{\prime}\rho^{\prime}}P_{\lambda^{\prime}\rho^{\prime}\mu\nu}\right\}\\ &-\mu^{3\epsilon}\frac{\Lambda\kappa^{4}\xi}{2}k_{1(\mu^{\prime}}k^{\prime}_{1\nu^{\prime})}P^{\mu^{\prime}\nu^{\prime}\alpha^{\prime}\beta^{\prime}}\eta_{\alpha^{\prime}\beta^{\prime}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}q^{2}}\left\{-\frac{1}{2}P_{\gamma\delta}{}^{\mu\alpha}P^{\gamma\delta}{}_{\mu\lambda}l^{\lambda}(l-q)_{\alpha}+\frac{3}{4}l\cdot(l-q)P_{\gamma\delta}{}^{\mu\nu}P^{\gamma\delta}{}_{\mu\nu}\right\}\\ =&\frac{\mu^{3\epsilon}m_{1}^{2}\Lambda\kappa^{4}\xi}{4}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{1}{l^{2}(l-q)^{2}}-\frac{\mu^{3\epsilon}\Lambda\kappa^{4}\xi}{q^{2}}\int\frac{d^{d}l}{(2\pi)^{d}}\frac{2k_{1\mu}k^{\prime}_{1\nu}l^{\mu}l^{\nu}+q^{2}l\cdot q/2}{l^{2}(l-q)^{2}}\end{split}$ (20)
 Only the first integral contributes to the potential at leading order, which is $\sim\ln q^{2}$.
-Likewise by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex.
-Putting things together now including 14, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
-$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}$ (19)
+Likewise, by replacing $m_{1}$ by $m_{2}$, we obtain the result if $k_{2}-k^{\prime}_{2}$ is the non-minimal two scalar-two graviton vertex.
+Putting things together now, including 16, we have the total Feynman amplitude for 2-2 scattering at ${\cal O}(\xi\Lambda\kappa^{4})$, in the non-relativistic limit,
+$-i{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}=-\frac{i\Lambda\kappa^{4}\xi m_{1}^{2}m_{2}^{2}}{2\pi^{2}\vec{q}^{2}}-\frac{i(m_{1}^{2}+m_{2}^{2})\Lambda\kappa^{4}\xi}{4\pi^{2}}\ln\frac{\vec{q}^{2}}{\mu^{2}}+\ {\cal O}\left(\vec{q}^{2}\ln\frac{\vec{q}^{2}}{\mu^{2}}\right).$ (21)
+Note that since $|\vec{q}|$ is non-relativistic, we must have $|\vec{q}|/m\ll 1$. We expect the contribution from the first term in the potential to dominate the other.
+Before we compute the potential corresponding to 21, a couple of comments on vanishing of some Feynman amplitudes are in order.
+Specifically, we have seen that two of the seagull diagrams falling into the topology of the first of 2 vanish.
+Likewise the amplitudes for all the vacuum polarisation diagrams, 3, vanish, whereas for 4, only two of the diagrams give non-vanishing contributions.
+Such vanishing amplitudes might correspond to some symmetry associated with the non-minimal and/or the $\Lambda$-three graviton vertices.
+For all these cases, note that the diagrams are nicely symmetric with respect to the position of the non-minimal vertex.
+We also note that two of the diagrams 4, even though they look symmetric with respect to the location of the non-minimal vertex, the said vertex is two scalar-two graviton and not two scalar-one graviton, and the diagrams yield non-vanishing contributions, 20.
+Thus the aforesaid symmetry, if there is any, seems to be only relevant to the non-minimal matter-graviton 3-point vertex, along with the $\Lambda$-three graviton vertex.
+In particular, from the structure of 17 it is not difficult to see that if we compute the one loop self energy of the graviton at $\Lambda\kappa^{2}$} due to one $\Lambda$-three graviton vertex and one usual three graviton vertex, the self energy would vanish.
+This corresponds to the cancellation between various terms arising due to the $\Lambda$-vertex.
+However, any specific nature or form of this speculated symmetry remains elusive to us so far.
+In this regard, we also refer our reader to e.g.
+[51] for demonstration of vanishing of one loop one and two point functions in a theory of inflation containing higher derivative terms for the inflaton.
+We also refer to [52] for vanishing of some scattering amplitudes off the Kerr black hole.
 4 The final result and discussion
@@
 $V(\vec{r})=-\frac{1}{4m_{1}m_{2}}\int\frac{d^{3}\vec{q}}{(2\pi)^{3}}e^{i\vec{q}\cdot\vec{r}}{\cal M}^{\rm tot.,non-analytic}|_{\rm NR}(\vec{q}^{2})$
-Using the above, we obtain from 19 the leading contribution of the non-minimal coupling parameter $\xi$ to the two body gravitational potential,
-$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{\mu l\hbar}{cr^{2}}\right],$ (20)
-where the factor $\mu l\hbar/c$ ensures dimensional consistency, with $l$ being a length scale.
-In the absence of a cosmological constant, there can be no contribution from $\xi$ to the potential, at least to this order of the perturbation theory.
-Note that while the right hand side of the momentum space expression 19 is perfectly dimensionally consistent, the Fourier transform of the logarithm necessitates the insertion of such factor.
-How do we fix $l$?
-For example, we may set $l=R_{\rm TA,max}$, i.e.
-the maximum turn around radius defined below 1, as one of the natural length scales in this context, beyond which this result cannot be trusted.
-However, this choice is not unique and anything concrete beyond this remains elusive to us so far.
-Note that such ambiguity is absent for the $\xi=0$ part of the potential, due to the appearance of different powers of $G$ in different terms, fixing such term uniquely [3, 4].
-(Here we do not have any term first order in $G$).
-Note that the second term in the potential diverges if either of the masses is vanishing.
-If we understand it correctly, we do not expect such behaviour from a non-relativistic quantity.
-This problem originates from the absence of the scalar rest mass term in the scalar-graviton non-minimal vertex, 6.
-Does this imply that the $\ln\vec{q}^{2}$ term of 19 should not be related to a physical gravitational potential, or, we simply accept 20 as it is?
-Once again, any clear reasoning in the favour of either of this remains elusive to us so far.
-Further note that the potential of 20, even though corresponds to pure short scale physics, contains two fundamental length scales (apart from $R_{\rm TA,max}$) — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$.
-We believe this is a novel feature encoded in a short scale physics computation.
-It seems an important task to further investigate the effect of $\xi$, as well as the vertex associated with $\Lambda$ on various perturbative quantum gravity phenomenon.
-As we have emphasised, this coupling is qualitatively different from the minimal ones.
-It would thus be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertices and RG perspectives.
-A look into the soft theorems might also be interesting in this context.
-We wish to come to these issues in our future publications.
+Using the above, we obtain from 21, 10, the leading contribution of the non-minimal coupling parameter $\xi$ to the two body long range gravitational potential,
+$V(G^{2},\Lambda,\xi,r)=-\frac{32G^{2}\Lambda\xi m_{1}m_{2}}{\pi r}\left[1-\frac{m_{1}^{2}+m_{2}^{2}}{m^{2}_{1}m^{2}_{2}}\frac{1}{r^{2}}\right]$ (22)
+We note the appearance of masses in the denominator of the second term on the right hand side.
+Each mass has to be understood as the inverse of the corresponding Compton wavelength.
+Thus the first term dominates the second.
+Perhaps a two loop computation for the class of diagrams that gave the first term would produce a result comparable to the second.
+Note also that this second term has originated from the Fourier transform of $\ln\vec{q}^{2}$, and any such logarithm originates from the term like $(q^{2}/\mu^{2})^{\epsilon/2}\cdot\epsilon^{-1}$ coming from 9.
+Hence it is essentially a quantum contribution, whereas the first is not.
+In this regard, we note the $\xi=0$ part of the gravitational potential up to one loop [3, 4],
+$V(\Lambda=0,r)|_{\rm minimal}=-\frac{Gm_{1}m_{2}}{r}\left[1-\frac{G(m_{1}+m_{2})}{r}-\frac{127G}{30\pi^{2}r^{2}}\right]$ (23)
+In 22, we do not have any term of first order in $G$ at all, due to the absence of tree level contribution.
+In fact the vanishing tree level long range contribution with $\xi\neq 0$ actually saves $\xi$, for otherwise a contribution like $\xi Gm_{1}m_{2}/r$ would have constrained $\xi$ very severely from simple solar system data, as is the case for the parameter of the Brans-Dicke theory [53].
+Second, we also note that one can use the $\Lambda$-3 graviton vertex used here to evaluate the two body potential with $\xi=0$.
+However, the result is expected to be subleading compared to the $\Lambda=0$ ones at our subhorizon scale.
+To the best of our knowledge, the vertex associated with $\Lambda$ has not been considered earlier.
+Note also that the potential of 22 contains two fundamental length scales — the shortest or Planck length corresponding to $G$, and the longest or Hubble horizon scale corresponding to $\Lambda$.
+We believe this is a novel feature encoded in physics of length scale much small compared to that of the horizon.
+Let us now compare the leading term of 22 with that of the ${\cal O}(G^{2}\xi)$ term with $\Lambda=0$ (${\cal O}(r^{-4})$) discussed below 12.
+The ratio of the first to the second for $m_{2}\gg m_{1}$ behaves as,
+$\sim\Lambda r^{2}\cdot\frac{m_{1}}{m_{2}}\cdot m_{1}r$
+Recall that a mass term is understood as its associated inverse Compton wavelength ($mc/h$).
+Imagine now that $m_{2},m_{1}$ are respectively the sun and the Mercury.
+Taking $\Lambda\sim 10^{-52}{\rm m}^{-2}$ and $r\sim 10^{10}{\rm m}$ to be the radius of the perihellion, the above ratio becomes $\sim 10^{36}$.
+It is clear that this ratio will keep increasing with increasing $r$ or $m_{1}$ or both.
+Taking now $m_{1}$ to be the electron, the ratio is instead found to be $10^{-72}$.
+Thus for ‘large’ $m_{1}$ values, the first term of 23 would dominate over the ${\cal O}(G^{2}\xi)$ terms.
+We believe the dominating role of a term associated with $\Lambda$ compared to the $r^{-4}$ term for large mass objects, even in the much subhorizon solar scale physics, is interesting in its own right.
+Let us now compare 22 with the standard Newton potential for two massive scalars.
+The leading parts of 22 and the above has a ratio of ${\cal O}(G\Lambda\xi)\sim\xi(L_{P}/L_{C})^{2}$, where $L_{C}\sim\Lambda^{-1/2}$ is the size of the cosmological event horizon and $L_{P}$ is the Planck length.
+Thus this ratio can at most be $10^{-10}$ in the context of primordial inflationary universe and much lower for our current universe.
+Also, the ratio of the leading part of 22 and the most subleading part of 23 reads $\sim\Lambda\xi r^{2}$, which is also small in the subhorizon scale we are interested in.
+We also note that the dependence on $m_{1}$ and $m_{2}$ for the two potentials, 22, 23, are qualitatively very different.
+This originates from the fact that 22 contains $\Lambda$, which has mass dimension two.
+Also as we have mentioned towards the end of 1, perturbative quantum gravity predictions in the flat background with a $\Lambda$ are expected to be subleading compared to that of the $\Lambda=0$ case, at least in our current universe where $\Lambda$ is very tiny.
+Such computations in the early inflationary universe might yield some even non-perturbative, phenomenologically significant contributions, e.g.
+[36].
+Note that in such background, computation of scattering at large scales is not possible due to vacuum instability.
+Instead the effective action formalism might be useful.
+We wish to come back to the issue of $\xi R\phi^{2}$ coupling and gravitational potential in the context of early inflationary universe in our future work.
+There are further things which seem to need our attention in the context of the matter-graviton non-minimal coupling.
+For example, a complete evaluation of the two body potential at various post-Minkowski order would be interesting.
+An investigation on the effect of $\xi$ on gravitational radiation seems to be an important task.
+These might lead to important constraint on $\xi$, giving us valuable insight about the matter-gravity interaction.
+It would also be interesting to investigate the effect of $\xi$ on self energy, various matter-graviton vertex functions, and RG perspectives.
+We wish to address these issues in our future publications.
 Acknowledgements : SB would like to thank Amitabha Lahiri and Parthasarathi Majumdar for useful discussions.
+The authors are indebted to anonymous referees for careful and critical reading of the manuscript and for making many useful comments and suggestions.